Will rate!!

Areser asnetehat shudets hat attand cssey mi close per m,geerally get better grades For the class, he verall percent of shufents who attnd regulerty in %Of those sho ceme to sna egular basi. ฐ9% rece ve As, of those who dont attend eegulahy. Only 10% get As, al students what propotion earm an A nd dant attend class relanly a nadent earned ษ。A, what → the chance fev อeend eins regularly?
Answer #1

a)P(R)=0.78

b)P(R')=1-0.78=0.22

c)P(A|R)=0.39

d)P(A'|R)=1-0.39=0.61

e)P(A|R')=0.1

f)P(A'|R')=1-0.1=0.9

b)

probability=P(R')*P(A|R')=0.22*0.1=0.022

c)

P(A)=P(R)*P(A|R)+P(R')*P(A|R')=0.78*0.39+0.22*0.1=0.326

d)

P(R|A)=P(R)*P(A|R)/P(A)=0.78*0.39/0.3262 =0.933